A 0.47 g sample of liquid C6H12 was combusted completely using excess oxygen inside a bomb (constant volume) calorimeter, with the products being carbon dioxide and liquid water. The calorimeter's heat capacity is 4.716 kJ °C-1.
If the temperature inside the calorimeter increased from 25.0 °C to 29.6 °C, determine ΔrH for this reaction in kJ mol-1 (with respect to C6H12) at 298 K.
1
Expert's answer
2019-10-23T07:01:08-0400

Moles of C6H12


"n(C_6H_{12})=\frac{m}{M} =\frac{0.47}{84.16} = 0.00558 mol"

Energy absorbed by calorimeter:


"q = c_{cal}\times \Delta T = 4.716\times (29.6-25)=21.69 kJ"


"\Delta_r H = -q"

"\Delta _r H = \frac{-q}{n} = \frac {-21.69 kJ}{0.00558 mol} =- 3887.74 \frac{kJ}{mol}"